OpenStudy (anonymous):

Please help! Solve: 2(x + 1) = 2x + 2 a. All Real Numbers b. No Solution c. 0 d. 1

OpenStudy (agentnao):

Distribute the 2 on the left side. Both sides are equal. You can insert ANY number and the equation will remain true. The answer is all real numbers.

OpenStudy (anonymous):

I would have got that wrong, thank you! Do you think you can help me with another one?

OpenStudy (agentnao):

Sure.

OpenStudy (anonymous):

Solve: -2(2x + 5) - 3 = -3(x - 1)

OpenStudy (agentnao):

Do you know what it means to distribute?

OpenStudy (agentnao):

Actually, have you learned the Order of Operations yet?

OpenStudy (anonymous):

Yes and yes. I tried by .. -4x - 10 - 3 = -3x + 3 Did I do that correctly?

OpenStudy (idkwut):

Yes.

OpenStudy (anonymous):

And then when it comes to choosing a number to take away, would I do .. the -3 on the left side to take out the other 3 on the right side. Which would give me -4x - 10 = -3x ?

OpenStudy (idkwut):

Yes, you subtract the 3 from -10, however, -10 - 3 does not equal -10.

OpenStudy (idkwut):

-4x-10-3 = -3x+3 -4x - 13 = -3x + 3 -4x = -3x + 16 -1x = 16 x = -16 Do you get it?

OpenStudy (anonymous):

Ohh okay, yes so you added the 13 and thats how you got 16. Thats the part I always get confused with but thank you!

OpenStudy (agentnao):

Oops! Sorry for not finishing the question with you! I got distracted. But I'm glad someone helped you! Let me know if you need help with anything else :)

OpenStudy (anonymous):

Lol that's okay. Alright and thank you also! (:

OpenStudy (idkwut):

No problem! :)